OpenStudy (anonymous):

PLEASE HELP I just need help with the last question. 1. Find me the midpoint between a and b given that A=(-1 5) and B=(7 -3) Answer - (3,1) 2. Calculate the distance between A and B. Answer - sqrt of 128 units 3.Find the equation in slope-intercept form of the line that is the perpendicular bisector of the segment between (-3,4) and (3,-8). Answer- I don't know

jimthompson5910 (jim_thompson5910):

What is the midpoint of (-3,4) and (3,-8)

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

2,-2

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

really?

jimthompson5910 (jim_thompson5910):

try again

jimthompson5910 (jim_thompson5910):

pay attention to the signs (negative/positive)

OpenStudy (anonymous):

\[\left[\begin{matrix}-3+3 ,&, 4+(-8) \\ 2 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

thats its right?

jimthompson5910 (jim_thompson5910):

-3+3 is 0 divide that by 2, you get 0/2 = 0

jimthompson5910 (jim_thompson5910):

so the midpoint should be (0,-2)

OpenStudy (anonymous):

woops I wasnt thinking strait I put 2

OpenStudy (anonymous):

what next

jimthompson5910 (jim_thompson5910):

now find the slope of the line through (-3,4) and (3,-8)

OpenStudy (anonymous):

how would i do that

jimthompson5910 (jim_thompson5910):

using the slope formula m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

-12/-1

OpenStudy (anonymous):

is that it? or am I completely wrong again?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

use this eq \[\frac{ y-y1 }{ x-x1 }=\frac{ y1-y2 }{ x1-x2 }\] solv it you will get the desired eq..!!!